OpenStudy (anonymous):

What percent of 30 is 240? My answer is 12.5. Can anyone check and see if thats right? :)

OpenStudy (radar):

12.5% is correct 30/240=.125 .125 x 100 = 12.5%

OpenStudy (anonymous):

Back to front....

OpenStudy (anonymous):

of..is not is...of

OpenStudy (radar):

Yeah I looked at it backwards

OpenStudy (radar):

It is 800%

OpenStudy (anonymous):

240/30 = 8 (In other words, it's 8 times bigger = 800%)

OpenStudy (radar):

240/30=8 8 X 100 = 800%

OpenStudy (anonymous):

30 = 100% 240 = 800% (multiply by 8)

myininaya (myininaya):

30n=240 n=240/30=24/3=8 we want a percentage though 8(1)=8(100%)=800%